OpenStudy (anonymous):

lim as x approaches 0+ (x^(-1/lnx))

OpenStudy (anonymous):

is this \[x^{-\frac{1}{ln(x)}}\]?

OpenStudy (anonymous):

in any case first step wil be to see this is of the form \[0^{-\infty}\] so take the log get \[-\frac{1}{\ln(x)}\ln(x)\] \[=-1\]

OpenStudy (anonymous):

so your answer will be \[e^{-1}=\frac{1}{e}\]